next up previous contents
Next: SOHO Intercalibration Committee Up: Minutes of the Previous: Status of the

Status of Experiments

In a special session held on Wednesday, June 15, the PI's presented the status of their experiments to DSci, R.Bonnet (for hardcopies of the PI presentations see annex 11). The session was opened by L.Culhane who gave a summary of the Task Force activities (cf. annex 10). After the PI presentations there was a visit to Matra Marconi where one could see the FM PLM. DSci expressed serious concerns on the viability of the SOHO mission if contamination of instruments was not better controlled by the PI's. In conclusion, DSci requested UVCS, CELIAS, EIT, and CDS PI's to report on 14 September 1994 at ESTEC on current technical and schedule status.

Bernhard Fleck
Tue Jan 6 15:25:49 EST 1998